# Variational Tail Bounds for Norms of Random Vectors and Matrices

### v3

@@ Appendix A A Tail bound for Euclidean norm of sub-exponential random vectors
${\left\lVert X\right\rVert}_{2}$ $\leq 4\sqrt{e}\eta\left(\sqrt{t\operatorname{tr}\left(\Sigma\right)}+t\sqrt{{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}}\right)\,.$
Proof.
Again we use Lemma˜1 with $\operatorname{\mathbb{P}}_{0}$ being the law of $U=\kappa_{0}G$ for a certain $\kappa_{0}>0$ and a standard Gaussian vector $G\in\mathbb{R}^{d}$. The assumption on $X$ provides the inequality
$M_{X}(u,2p)$ $\leq\eta^{2p}(2p)^{2p}{\left\lVert u\right\rVert}^{2p}_{\Sigma}\,$,
which together with Lemma˜3 yields
$\left(\operatorname{\mathbb{E}}_{0}\left(M_{X}(U,2p)\right)\right)^{1/(2p)}$ $\leq 2\eta p\kappa_{0}\left(\operatorname{\mathbb{E}}\left({\left\lVert G\right\rVert}_{\Sigma}^{2p}\right)\right)^{1/(2p)}$
@@
$\leq 2\eta p\kappa_{0}\left(\operatorname{tr}(\Sigma)+p{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}\right)^{1/2}\,.$
Furthermore, using (11) we have
$\operatorname{\mathbb{P}}_{0}\left(\left|{\langle U,x\rangle}\right|\geq{\left\lVert x\right\rVert}_{2}\right)$ $=2\Phi\left(-\frac{1}{\kappa_{0}}\right)\geq e^{-2/(3\kappa_{0}^{2})}/2\,$,
and thus
It then follows from Lemma˜1 (considering the moment of order $2p$ rather than $p$) that with probability at least $1-e^{-t}$ we have
${\left\lVert X\right\rVert}_{2}$ $\leq\inf_{\begin{subarray}{c}p\geq 1\\ \kappa_{0}>0\end{subarray}}2^{1+1/(2p)}\eta p\kappa_{0}e^{1/(3p\kappa_{0}^{2})}\left(\operatorname{tr}\left(\Sigma\right)+p{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}\right)^{1/2}e^{t/p}$
$\leq\inf_{p\geq 1}4\eta\left(p\operatorname{tr}\left(\Sigma\right)+p^{2}{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}\right)^{1/2}e^{t/(2p)}$
@@
$\leq 4\sqrt{e}\eta\left(\sqrt{t\operatorname{tr}\left(\Sigma\right)}+t\sqrt{{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}}\right)\,.$
where the last line follows by choosing $p=t$ and using subadditivity of the square root. ∎
@@ Appendix B Supporting Lemmas
We can view (13) as a projection (in a Hilbert space with norm ${\left\lVert\cdot\right\rVert}_{\Sigma^{1/2}}$). Therefore, using the fact that projections onto closed convex sets are firmly nonexpansive [BC17, Propostion 4.8] and thus define contractions, for all $x,y\in\mathbb{R}^{d}$ we have

### v4

@@ Appendix A Recovering Example˜2 and (13) via Theorem˜2
$\Omega_{\Sigma}(p,\kappa_{0},\infty)$ $=\eta\sqrt{p}\kappa_{0}\left(\sqrt{\operatorname{tr}(\Sigma)}+\sqrt{2p{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}}\right)\,.$
$=\frac{\kappa_{0}}{\sqrt{2\pi}}e^{-\kappa_{0}^{-2}}\,$
Therefore, the bound (15) of Theorem˜2 implies
@@
${\left\lVert X\right\rVert}={\left\lVert X\right\rVert}_{2}$ $\leq\inf_{p\geq 1,\,\kappa_{0}>0}e^{t/(2p)}\eta\sqrt{p}\kappa_{0}\left(\sqrt{\operatorname{tr}(\Sigma)}+\sqrt{2p{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}}\right)8^{1/(2p)}e^{1/(2p\kappa_{0}^{2})}$
@@
$=\sqrt{e}\eta\inf_{p\geq 1}e^{(t+\log 8)/(2p)}\left(\sqrt{\operatorname{tr}(\Sigma)}+\sqrt{2p{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}}\right)$
$\leq 4\sqrt{e}\eta\left(3\operatorname{tr}(\Sigma)+\frac{2}{3\log 2}{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}t\right)^{1/2}\,$,
where the third line follows by evaluating the argument of the infimum at $p=1+t/\log 8$, the Cauchy–Schwarz inequality, and the fact that ${\left\lVert\Sigma\right\rVert}_{\mathrm{op}}\leq\operatorname{tr}(\Sigma)$. Therefore, up to the constant factors, we have recovered the bound (13) through Theorem˜2.
@@ Appendix B Supporting Lemmas
We can view (16) as a projection (in a Hilbert space with norm ${\left\lVert\cdot\right\rVert}_{\Sigma^{1/2}}$). Therefore, using the fact that projections onto closed convex sets are firmly nonexpansive [12, Propostion 4.8] and thus define contractions, for all $x,y\in\mathbb{R}^{d}$ we have